Fermi National Laboratory


Accelerator Update

Wednesday January 7
The day shift began with store #3130 colliding, stacking in the Antiproton Source (Pbar), and with beam to MiniBoone.

At 1:08 PM the anode supply for Pbar RF station 1 tripped off. The power supply had to be replaced, stacking resumed at 2:12 PM.

Tevatron (TeV) studies began at 6:10 PM, with the store being terminated at 8:25 PM.

Shot setup began at 10:13 PM.

At 11:43 PM two power supplies in Main Injector (MI) were tripped off. Supplies would not reset and an expert was called in to investigate. Power supplies were back on at 12:44 AM.

Thursday January 8
The midnight shift began with the Tev in shot setup and with MiniBoone taking beam.

At 1:57 AM, Operations established store #3132 with an initial luminosity of 49.11 E30.

Pbar returns to stacking at 2:01 AM.

At 7:32 AM shot setup for Pbars to the Recycler Ring (RR) began, and was completed by 9:39 AM.

Various studies were performed for SY120, slip stacking, and recycler.

Friday January 9
The Midnight shift began with store #3132 colliding, stacking, Sy120 studies, and with beam to MiniBoone.

TeV studies began at 7:30 AM, with the store being terminated at 8:18 AM. A Tev quench occurred at F4 while the beam terminated.

Store #3132 set a new one store luminosity record of 2.546 pb-1.

Today's Plans
Recover from the Tev quench, followed by Tev studies, then go into shot setup.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 1/9/2004   email Fermilab